OpenStudy (anonymous):

help...what am i doing wrong? The function f(x)=lnx has a Taylor series at a=9. Find the first 4 nonzero terms in the series, that is write down the Taylor polynomial with 4 nonzero terms. i tried ln(9)+1/9-1/9^2+2/9^3 and that's not working...

OpenStudy (koikkara):

forgot all these..sorry..

OpenStudy (phi):

see https://en.wikipedia.org/wiki/Taylor_series#Definition I think you left out the (x-9)^n terms

OpenStudy (phi):

and the divide by n! ? the first 3 terms are \[\ln(9) + \frac{(x-9) }{9} -\frac{ (x-9)^2 }{2\cdot 9^2}\]

OpenStudy (anonymous):

ok, let me try that. i've seen so many different formulas for stuff..

OpenStudy (anonymous):

ok. thanks, @phi. that worked. i was using the wrong formla